OpenStudy (anonymous):

Which number correctly completes this equation? 6 --- % * 45 = _____ 10 Select one of the options below as your answer: A. .027 B. .27 C. 2.7 D. 27

OpenStudy (inkyvoyd):

How much is 6/10 expressed in a decimal?

OpenStudy (anonymous):

To find a percent you take the number given (6 tenths) and divide by 100, Then multiply by 45.

OpenStudy (anonymous):

60%

OpenStudy (anonymous):

No. In no way should you get 60% in this equation. Your new equation to account for the percent is \[((6/10)/100) \times 45\]

OpenStudy (inkyvoyd):

I knew you would confuse him :P

OpenStudy (inkyvoyd):

This is a trick question, @michaelJordan

OpenStudy (inkyvoyd):

6/10% is not 6/10

OpenStudy (anonymous):

It's not a trick question if you read % as /100.

OpenStudy (inkyvoyd):

6/10%=(0.6)*0.01 6/10=0.6

OpenStudy (inkyvoyd):

It is a trick question to most people learning percents.

OpenStudy (anonymous):

That's why I was trying to explain that a percent is the number given before the % sign divided by 100.

OpenStudy (anonymous):

how do i get the answer?